OpenStudy (anonymous):

\[\sqrt{8}-\sqrt{2}\] how do you get sqrt 2 as your answer

OpenStudy (agreene):

\[\sqrt{8}-\sqrt{2}\] \[2\sqrt{2}-\sqrt{2}\] \[(\sqrt{2}*\sqrt{2})-\sqrt{2}\] \[(\sqrt{2}*\cancel{\sqrt{2}})-\cancel{\sqrt{2}}\] \[\sqrt{2}\]

OpenStudy (anonymous):

\[\sqrt{8}= \sqrt{4\times2}=2\sqrt{2}\] \[2\sqrt{2}-\sqrt{2}= \sqrt{2}\]

OpenStudy (anonymous):

After you guys are done with this question, can someone help me with mine?

OpenStudy (anonymous):

solution given by agreene is incorrect